# Learning to Compress Contexts for Efficient Knowledge-based Visual Question Answering

### v1

@@ Methods
The lengths of the two sets of learnable prompts are hyperparameters. We conducted a series of comparative experiments on these two hyperparameters, with the results presented in subsequent sections. In most of our experiments, we set $L(\theta_{vq})$ to 12 and $L(\theta_{d})$ to 16. The initialization weights of learnable prompts also play a crucial role in the hyperMLLM’s ability to compress input contexts, particularly in the early training stage. We propose a strategy for initializing learnable prompts called PIPE, i.e. Prompt Initialization with hard Prompt Embeddings. We begin by manually designing two sets of hard prompts. For example, the hard prompt corresponding to $\theta_{d}$ is “Summarize the key information of the given passage in a concise manner.” The hard prompts are then processed by the tokenizer and embedding layer of the hyperMLLM to generate their embeddings. These embeddings are subsequently used to initialize the weights of the learnable prompts $\theta_{d}$ and $\theta_{vq}$.
Learning to aggregate the compressed contexts
Now given the set of compressed prompts of the retrieved documents $\{\theta_{i}\}^{K}$, $\theta_{v_{i}}$, $\theta_{q_{i}}$ and $\theta_{vq_{i}}$, RACC aims to use $\theta_{vq_{i}}$ as a query to aggregate more relevant information from $\{\theta_{i}\}^{K}$. Considering the characteristics of KB-VQA, we identify three key problems in designing the aggregator network and propose three corresponding strategies to address them.
How to grasp the relationship between V and Q in VQA? We first explore the relationship between V (images) and Q (questions) from the task and model perspectives.
From the perspective of KB-VQA, both images and questions are crucial components and are complementary to each other. However, when it comes to a specific image-question pair, the importance of the image and question may differ. In some cases, it is necessary to retrieve relevant documents based on key details present in the image but not mentioned in the question. Conversely, when the question contains substantial information, it is more appropriate to focus on retrieving documents closely related to the question. Given that the retrieved documents contain a lot of tokens and the portion of content relevant to the image or question might be relatively small, we believe that enhancing the semantic information in retrieved documents using image and question-based information is essential.
From the perspective of MLLMs, the relationship between images and questions is unequal, which is reflected in the number of tokens they occupy. In nearly all MLLMs, image features are converted into fixed-length tokens, such as 32, while the number of tokens for questions is often smaller. This disparity may result in the learnable prompts focusing too much on the image part in the early training stage.
Based on the above discussion, we propose a strategy called Decoupled Compression of Vision and Question (DCVQ). Our strategy begins by decoupling the given image-question pair $(v_{i},q_{i})$ and input them separately into the hyperMLLM with $\theta_{vq}$, generating $\theta_{v_{i}}$ and $\theta_{q_{i}}$, which respectively represent the compressed prompts of $v_{i}$ and $q_{i}$. At the same time, both $v_{i}$ and $q_{i}$ are used to retrieve $K$ documents from the knowledge source, which are then compressed into $\{\theta_{i}\}^{K}$. Then, we concatenate the obtained $\theta_{v_{i}}$ and $\theta_{q_{i}}$ and pass them through a cross-attention block with $\{\theta_{i}\}^{K}$ as the query. Through cross-attention computations, the semantic information in $\{\theta_{i}\}^{K}$ that is more relevant to $\theta_{v_{i}}$ as well as $\theta_{q_{i}}$ can be directly enhanced. This strategy also helps prevent learnable prompts from overly focusing on image tokens during the process of learning coupled and decoupled compression of image-question pairs. Denoting a naive cross-attention block (Vaswani et al. 2017) as CA, the computational process of DCVQ is represented as follows:
@@
where $\{\theta^{*}_{i}\}^{K}$ denotes the query-enhanced compressed prompts of retrieved documents.
@@
How to utilize the document retrieval scores to guide the aggregation process? During the retrieval process of most retrievers, they assign a confidence score to each retrieved document based on metrics such as embedding similarity. In the following text, we will use the term ”retrieval score” to represent this score. While determining whether a document can truly answer a question based on an image remains challenging, the retrieval score offers a relatively reliable metric for this purpose. Based on the principles of retrieval mechanisms, documents with higher retrieval scores are generally considered to provide more relevant and useful information for the given image and question. Lin et al. (2024a) uses the retrieval scores as a reference metric for selecting the final answer during the inference process, but it does not utilize this crucial metric in the training process. We try to utilize the retrieval scores to guide the process of aggregating information from multiple documents’ compressed prompts. Based on the above, we propose a strategy called Retrieval-Guided Cross-Attention (RGCA), which integrates retrieval confidence scores into the attention computation process of the original cross-attention mechanism.
@@
The retrieval-guided cross-attention block not only considers the embedding similarity between compressed prompts but also assigns more attention to the compressed prompts corresponding to documents with higher retrieval scores. We denote the retrieval-guided cross-attention block as $\textbf{CA}_{r}$, a forward pass of this module can be expressed as:
$\theta^{*}_{vq_{i}}=\textbf{CA}_{r}(\theta_{vq_{i}},\{d^{*}_{i}\}^{K},\{p_{i}\}^{K})$ (4)
@@
The number of retrieval-guided cross-attention blocks contained in the aggregator of our framework is set to $n_{r}=3$.
How to deal with the irrelevant documents? When using document bases as the external knowledge source for KB-VQA, the documents retrieved by the multimodal retriever may sometimes be completely irrelevant. Even if they are relevant, they may not provide the model with useful information to give the correct answer. Luo et al. (2021); Lin et al. (2024a) consider a document to be pseudo-relevant if it contains any of the human-annotated answers. RAVQA-v2 does not consider the impact of irrelevant documents and treats both irrelevant documents and useful documents in the same way, i.e. inputting them with corresponding image-question pairs into MLLMs for loss calculation. RAVQA-v2 forces MLLMs to generate correct answers even based on irrelevant documents, which imposes incorrect supervised signals on the MLLM and may cause it to generate information that is absent from the documents after finetuning.
Therefore, to better tackle the effects of irrelevant retrieved documents, we propose a strategy called Pseudo-Relevance-based Backpropagation Dropout (PRDB), which only computes the gradients of the corresponding compressed prompts of the pseudo-relevant documents.
Specifically, after hyperMLLM converts all documents into compressed prompts, we apply a stop gradient operation, i.e. stopgrad($g_{hyper}({d^{k}_{i}},\theta_{d})$) to the compressed prompts of those irrelevant documents before they are passed into the aggregator. In this way, after loss calculation and gradient backpropagation, the gradients of the compressed prompts corresponding to irrelevant documents will be truncated, preventing them from leading the learnable prompts’ weights to update in the wrong direction.
Modulating the baseMLLM with aggregated prompts
@@
After aggregating the compressed prompts in the aggregator network, we have the documents-based compressed prompts of vision and question, which is denoted as $\theta^{*}_{vq_{i}}$. Then we convert $\theta^{*}_{vq_{i}}$ into a P-Tuning v2 modulation for the downstream baseMLLM, which involves adding a small amount of KV cache at each layer of the baseMLLM. Given that different layers of MLLM process information at varying levels of abstraction and complexity, we employ a set of $m$ Multi-Layer Perceptrons (MLPs) for projecting $\theta^{*}_{vq_{i}}$ into additional KV cache of each layer in the baseMLLM. $m$ is the number of layers in the baseMLLM. Here, we denote the P-Tuning v2 modulation as $\Theta_{i}$ and the baseMLLM as $g_{base}$. Our framework can be optimized in an end-to-end manner using the loss function $\mathcal{L}$, namely the language modeling loss based on the ground truth answer:
$\underset{\theta_{d},\theta_{vq},h}{\min}\frac{1}{N}\sum^{N}_{i=1}\mathcal{L}(g_{base}(v_{i},q_{i};\Theta_{i}),a_{i})$ (5)
@@
where $h$ includes a single CA block, $n_{r}$ $\textbf{CA}_{r}$ blocks and a set of MLPs. $N$ is the batch size of training inputs. Depending on whether the hyperMLLM and baseMLLM are initialized from the same MLLM, our framework offers two variants: RACC-homogeneous and RACC-heterogeneous, abbreviated as RACC-homo and RACC-hetero, respectively.
Experiments
Datasets and Knowledge Sources
@@
We evaluate our framework on OK-VQA (Marino et al. 2019), which is the most widely studied KB-VQA dataset. We also conducted experiments on AOK-VQA (Schwenk et al. 2022), which is the successor of OK-VQA.
@@
In terms of the knowledge source, following Lin et al. (2024a), we adopt Google Search (Luo et al. 2021) for OKVQA and AOKVQA, which is a textual document base comprised of nearly 200 thousand documents. We also carefully curated a multimodal document source from the Wikipedia Image-Text dataset (Srinivasan et al. 2021) for OK-VQA to further demonstrate the applicability of our framework on multimodal knowledge sources. In this paper, we use GS and WIT to refer to these two knowledge sources.
@@
We use FLMR (Lin et al. 2024a) and PREFLMR (Lin et al. 2024b) as retrievers for document retrieval. The FLMR retriever was used to retrieve information from the GS knowledge source, while the PREFLMR retriever was used to retrieve information from the WIT knowledge source.
Training Setup
@@
Evaluation
We evaluate the performance of our framework using the official VQA Accuracy (Marino et al. 2019). Let $a_{i}$ be the list of human-annotated answers of the given image-question pair $(v_{i},q_{i})$, and $y_{i}$ be the model’s outputs. The VQA accuracy for $(v_{i},q_{i})$ is calculated as follows:
@@
Comparative Study
In this section, we will elaborate on the advantages of RACC compared to previous works from three aspects: performance, cost, and inference efficiency.
First of all, RACC outperforms many competitive baselines. The performance comparison of RACC and other competitive baselines on the OK-VQA dataset is presented in Table 1. Based on InstructBLIP-FlanT5XL, RACC-homo with GS as the knowledge source reaches an accuracy of 59.65%. With WIT as the knowledge source, our framework achieves 59.17%. When adopting RACC-hetero, with InstructBLIP-FlanT5XL as the hyperMLLM and InstructBLIP-Vicuna7B as the baseMLLM, we achieve a state-of-the-art (SOTA) accuracy of 62.9%.
@@
Table 3: Comparison of inference efficiency between RAVQA-v2 and RACC when adopting the WIT knowledge source. “Eval time” and “Disk Space” are metrics measured for a single image-question pair input, while “VQA Accuracy” represents the performance on the OK-VQA test set. “w pre” indicates that pre-saving the compressed prompts of retrieved documents before inference. The MLLM used in both two frameworks is InstructBLIP-FlanT5XL.
@@
In terms of cost, our work has notable advantages. First, we do not utilize any image-based textual descriptions provided by external APIs or models (Gui et al. 2021; Lin and Byrne 2022; An et al. 2024), such as captions, object tags, OCR etc. Second, RACC does not use any very large LLMs (ChatGPT, GPT-3) or MLLMs (GPT-4) but still achieves excellent performance even with small-scale MLLMs.
@@
The inference efficiency is the main concern of this paper. RACC demonstrates significant advantages in inference efficiency compared to RAVQA-v2 (Lin et al. 2024a). RACC not only significantly reduces inference latency but also minimizes disk space usage by pre-saving compressed prompts corresponding to the documents of the knowledge source. We present a comparison of the inference efficiency between RACC and RAVQA-v2 (Lin et al. 2024a) in Table 3. When pre-saving compressed prompts, we achieve a substantial reduction of 59.7% in inference latency and 91.0% in disk space usage. Even without pre-saved compressed prompts, the inference latency can still be reduced by 22.0%.
Ablation Studies
@@
We propose four strategies to improve the aggregation process of compressed contexts and conduct ablation studies to verify their effectiveness. The settings and results of ablation studies are depicted in Table 4 and Table 5. Note that we adopt RACC-homo with the GS knowledge source in the ablation studies, where the hyperMLLM and baseMLLM are both initialized from InstructBLIP-FlanT5XL.
@@
Firstly, comparing lines 2 and 3, as well as lines 7 and 8, we can observe that the PIPE strategy brings improvements of 0.31% and 0.42% under different settings. From the difference between lines 3 and 4 in Table 4, we observe that the DCVQ strategy brings an improvement of 0.77%. On the other hand, the RGCA strategy results in a performance gain of 0.37%, as shown in lines 3 and 5. Last but not least, the performance difference between lines 6 and 8 shows that the PRDB strategy leads to a performance gain of 0.54%.
@@
We also explore how to set the length of learnable prompts (i.e. $L(\theta_{vq})$ and $L(\theta_{d})$), and the results are shown in Table 5. We select the best configuration, setting $L(\theta_{vq})$ and $L(\theta_{d})$ to 12 and 16, respectively. All other experiments in this paper are conducted using this configuration. In the supplementary materials, we provide additional results of ablation studies and comparative experiments on the hyperparameter $K$.
No. PIPE DCVQ RGCA PRDB VQA Accuracy (%)
@@
Table 4: The results of ablation studies on the design of our aggregator module. The GS knowledge source is adopted here. The ablation experiments are conducted based on RACC-homo with InstructBLIP-FlanT5XL.
@@
Table 5: RACC-homo’s results of the comparative experiments on the length of the predefined learnable prompts $\theta_{vq}$ and $\theta_{d}$. “w/o PIPE” means that the learnable prompts are randomly initialized here.
Broad Applicability of RACC

### v2

@@ 3.2 Phase 1: Compression Learning
Solution: Therefore, we propose a method to initialize learnable prompts called Prompt Initialization with hard Prompt Embeddings (PIPE). We begin by manually designing two sets of hard prompts. For example, the hard prompt corresponding to $\theta_{d}$ is “Summarize the key information of the given passage in a concise manner.” The hard prompts are then processed by the tokenizer and the embedding layer of hyperMLLM to generate their embeddings. Subsequently, these embeddings are used to initialize the weights of the learnable prompts $\theta_{d}$ and $\theta_{vq}$.
Issue 2: How to deal with the irrelevant documents?
Analysis: The documents retrieved by the multimodal retriever may sometimes be completely irrelevant to the given image and question. Even if they are relevant, they may not provide the model with useful information to give the correct answer.
However, RAVQA-v2 (Lin et al., 2024a) does not consider the impact of irrelevant documents and treats both irrelevant documents and useful documents in the same way, i.e. concatenating retrieved documents with corresponding image-question pairs and inputting them into MLLMs to calculate the loss based on the correct answer. In clearer terms, RAVQA-v2 forces MLLMs to generate correct answers even based on irrelevant documents, which imposes incorrect supervised signals on the MLLM and can significantly harm the training process.
Solution: Therefore, to avoid the negative impact on the compression learning process caused by the involvement of irrelevant documents in training, we propose a method called Pseudo-Relevance-based Backpropagation Dropout (PRDB), which is introduced as follows:
Following Luo et al. (2021); Lin et al. (2024a), we consider a document to be pseudo-relevant if it contains any of the human-annotated answers. After all $K$ retrieved documents are converted into compressed prompts, we apply a stop gradient operation, i.e. stopgrad($\textsc{M}_{hyper}({d^{k}_{i}},\theta_{d})$) to the compressed prompts of those documents which are considered pseudo-irrelevant. Only the gradients of the compressed prompts of the pseudo-relevant documents will be backpropagated to the learnable prompts. In this way, after loss calculation and gradient backpropagation, PRDB prevents the gradients of the compressed prompts corresponding to irrelevant documents from leading the learnable prompts’ weights to update in the wrong direction, ensuring the stability of the compression learning phase.
3.3 Phase 2: Information Aggregation
After the compression learning phase, we now have a set of compressed prompts of the retrieved documents $\{\theta_{i}\}^{K}$, $\theta_{v_{i}}$, $\theta_{q_{i}}$ and $\theta_{vq_{i}}$. In this section, we detail the process of using $\theta_{vq_{i}}$ to aggregate useful semantic information from $\{\theta_{i}\}^{K}$. We further identify two key issues in designing the information aggregation process and propose two corresponding methods to address them. We name this phase as the information aggregation phase.
Issue 3: How to enhance the semantic information in the compressed prompts of documents associated with the given image and question?
Analysis: The retrieved documents often contain a large number of redundant tokens, with only a small portion of tokens being relevant to the question and the image, which can be useful for answering the question and should be paid more attention to. Even after compression, the compressed prompts can still contain redundant semantic information. We think that leveraging the information from the image and the question to enhance the semantic information in the compressed prompts of documents is essential.
Solution: Based on the above, we propose a method called DCSE, i.e. Decoupled Compression for Semantics Enhancement. Our method begins by decoupling the given image-question pair $(v_{i},q_{i})$ and input them separately into the hyperMLLM with $\theta_{vq}$, generating $\theta_{v_{i}}$ and $\theta_{q_{i}}$, which can be denoted as follows:
$\begin{split}&\theta_{v_{i}}=\textsc{M}_{hyper}(v_{i},\theta_{vq})\\ &\theta_{q_{i}}=\textsc{M}_{hyper}(q_{i},\theta_{vq})\end{split}$ (3)
Compared with $\theta_{vq_{i}}$ obtained by jointly compressing the image-question pair, $\theta_{v_{i}}$ and $\theta_{q_{i}}$ derived from decoupled compression can better preserve the semantic information contained in both the image and the question. In the retrieved documents, the semantic information related to either the image or the question is of significance for answering the question. Therefore, we employ the concatenated $\theta_{v_{i}}$ and $\theta_{q_{i}}$ to enhance the semantic information of $\{\theta_{i}\}^{K}$ via cross-attention. Here, we adopt a original cross-attention block, where $\{\theta_{i}\}^{K}$ serves as the query, and the concatenated $\theta_{v_{i}}$ and $\theta_{q_{i}}$ act as both the key and the value. Denoting a original cross-attention block (Vaswani et al., 2017) as CA, the computational process of DCSE is denoted as follows:
@@ 3.3 Phase 2: Information Aggregation
$\{\theta^{*}_{i}\}^{K}$ represents the query-enhanced compressed prompts of documents.
Issue 4: How to utilize the document retrieval scores to guide the aggregation process?
Analysis: In the retrieval process, most retrievers assign a confidence score to each retrieved document based on metrics such as embedding similarity. In the following text, we use ”document retrieval score” to represent this score.
@@
While judging whether a document can truly answer a question based on an image is difficult, the retrieval score offers a reliable metric for this purpose. Based on the principles of retrieval mechanisms, documents with higher retrieval scores are generally considered to provide more relevant and useful information for the given image and question. Lin et al. (2024a) utilizes the retrieval scores as a reference metric for selecting the final answer in the inference process, but it does not use this crucial metric in the training process.
@@
Solution: Based on the above, we propose a method called Retrieval-Guided Cross-Attention (RGCA). RGCA is an improvement of the original cross-attention mechanism, designed to gather the semantic information from $\{\theta^{*}_{i}\}^{K}$ that can assist in answering questions into $\theta_{vq_{i}}$, guided by $\{p_{i}\}^{K}$. RGCA not only considers the embedding similarity between compressed prompts but also assigns more attention to the compressed prompts of documents with higher retrieval scores. Refer to the appendix for the pseudo code of RGCA. We denote a retrieval-guided cross-attention block as $\textbf{CA}_{r}$, whose forward pass is denoted as:
$\theta^{*}_{vq_{i}}=\textbf{CA}_{r}(\theta_{vq_{i}},\{\theta_{i}^{*}\}^{K},\{p_{i}\}^{K})$ (5)
@@
The number of retrieval-guided cross-attention blocks contained in our framework is set to $n_{r}=3$.
3.4 Phase 3: Modulation Generation
@@
After the information aggregation phase, we have the documents-based compressed prompts of vision and question, which is denoted as $\theta^{*}_{vq_{i}}$. Then we convert $\theta^{*}_{vq_{i}}$ into a P-Tuning v2 modulation for the downstream baseMLLM, which involves adding a small amount of KV cache at each layer of the baseMLLM. Given that different layers of MLLM process information at varying levels of abstraction and complexity, we employ a set of $m$ Multi-Layer Perceptrons (MLPs) for projecting $\theta^{*}_{vq_{i}}$ into the additional KV cache of each layer in the baseMLLM, where $m$ is the number of layers within the baseMLLM. Here, we denote the generated P-Tuning v2 modulation as $\Theta_{i}$ and the frozen baseMLLM as $\textsc{M}_{base}$.
Our framework can be optimized in an end-to-end manner using the loss function $\mathcal{L}$, namely the language modeling loss based on the ground truth answer:
$\underset{\theta_{d},\theta_{vq},h}{\min}\frac{1}{N}\sum^{N}_{i=1}\mathcal{L}(\textsc{M}_{base}(v_{i},q_{i};\Theta_{i}),a_{i})$ (6)
@@
$h$ includes a single CA block, $n_{r}$ $\textbf{CA}_{r}$ blocks and a set of MLPs. $N$ is the training batch size. RACC offerss two variants: RACC-homogeneous and RACC-heterogeneous, abbreviated as RACC-homo and RACC-hetero, respectively. In the setup of RACC-homo, the hyperMLLM and the baseMLLM are identical, which means that the MLLM learns to compress contexts for itself. For RACC-hetero, the hyperMLLM and the baseMLLM differ in either structure or weight.
4 Experiments
4.1 Datasets and Knowledge Sources
@@
We evaluate our framework on OK-VQA (Marino et al., 2019), which is the most widely studied KB-VQA dataset. We also conduct experiments on AOK-VQA (Schwenk et al., 2022), which is the successor of OK-VQA.
@@
In terms of the knowledge source, following Lin et al. (2024a), we adopt Google Search (Luo et al., 2021) for OK-VQA and AOK-VQA, which is a textual document base comprised of nearly 200 thousand documents. We also carefully curated a multimodal document source from the Wikipedia Image-Text dataset (Srinivasan et al., 2021) for OK-VQA to show that our framework also works well with multimodal knowledge sources. We use GS and WIT to refer to these two knowledge sources.
@@
We use FLMR (Lin et al., 2024a) and PREFLMR (Lin et al., 2024b) as retrievers for document retrieval. The FLMR retriever was used to retrieve information from the GS knowledge source, while the PREFLMR retriever was used to retrieve information from the WIT knowledge source.
4.2 Training Setup
@@
4.3 Evaluation
We evaluate the performance of our framework using the official VQA Accuracy (Marino et al., 2019). Let $a_{i}$ be the list of human-annotated answers of the given image-question pair $(v_{i},q_{i})$, and $y_{i}$ be the model’s outputs. The VQA accuracy for $(v_{i},q_{i})$ is calculated as follows:
@@ 4.3 Evaluation
4.4 Comparative Study
In this section, we will elaborate on the advantages of RACC compared to previous works from the following three aspects: performance, cost, and inference efficiency.
First of all, RACC outperforms many competitive baselines. The performance comparison of RACC with other competitive baselines on the OK-VQA dataset is presented in Table 1. Based on InstructBLIP-FlanT5XL, RACC-homo with GS as the knowledge source reaches an accuracy of 59.65%. With WIT as the knowledge source, our framework achieves 59.17%. When adopting RACC-hetero, with InstructBLIP-FlanT5XL as the hyperMLLM and InstructBLIP-Vicuna7B as the baseMLLM, we achieve a state-of-the-art (SOTA) accuracy of 63.92%.
@@ 4.4 Comparative Study
Table 3: Comparison of inference efficiency between RAVQA-v2 and RACC when adopting the WIT knowledge source. “Eval time” and “Disk Usage” are measured for a single image-question pair input. “w pre” indicates pre-saving the compressed prompts of retrieved documents before inference. The MLLM used in both two frameworks is InstructBLIP-FlanT5XL.
@@
In terms of cost, our work has notable advantages. First, we do not utilize any image-based textual descriptions provided by external APIs or models (Gui et al., 2021; Lin & Byrne, 2022; An et al., 2024), such as captions, object tags, OCR etc. Second, RACC does not use any very large LLMs (ChatGPT, GPT-3) or MLLMs (GPT-4) but still achieves excellent performance even with small-scale MLLMs.
@@
The inference efficiency is the main concern of this paper. RACC demonstrates significant advantages in inference efficiency compared to RAVQA-v2, which is shown in Table 3. In Table 3, the inference efficiency and disk usage of RACC are demonstrated under two scenarios: ”w/o pre” and ”w pre”. Here, ”w pre” refers to the scenario where the compressed encodings of retrieved documents are pre-saved before the inference process, which is shown in Figure 1. We also present the inference efficiency and disk usage of RAVQA-v2. When pre-saving compressed prompts, we achieve a substantial reduction of 59.7% in inference latency and 91.0% in disk space usage compared to RAVQA-v2. Even without pre-saved compressed prompts, the inference latency can still be reduced by 22.0%.
4.5 Ablation Studies
@@
We propose four methods to improve the aggregation process of compressed contexts and conduct ablation studies to verify their effectiveness. The settings and results of ablation studies are depicted in Table 8 and Table 5. Note that we adopt RACC-homo with the GS knowledge source in the ablation studies, where the hyperMLLM and baseMLLM are both initialized from InstructBLIP-FlanT5XL. Refer to the appendix for more details of the ablation studies.
@@
Firstly, comparing lines 2 and 3, as well as lines 7 and 8, we can observe that the PIPE method brings improvements of 0.31% and 0.42% under different settings. From the difference between lines 3 and 4 in Table 8, we observe that the DCSE method brings an improvement of 0.77%. On the other hand, the RGCA method results in a performance gain of 0.37%, as shown in lines 3 and 5. Last but not least, the performance difference between lines 6 and 8 shows that the PRDB method leads to a performance gain of 0.54%.
@@
We also investigate into the settings of some important hyperparameters. We first explore how to set the length of learnable prompts (i.e. $L(\theta_{vq})$ and $L(\theta_{d})$), and the results are shown in Table 5. We select the best configuration, setting $L(\theta_{vq})$ and $L(\theta_{d})$ to 12 and 16, respectively. All other experiments in this paper are conducted using this configuration. In the appendix, we provide comparative experiments on the hyperparameter $K$.
No. PIPE DCSE RGCA PRDB VQA Accuracy (%)
@@
Table 4: The results of ablation studies of RACC. The GS knowledge source is adopted here. The ablation experiments are conducted based on RACC-homo with InstructBLIP-FlanT5XL.
@@
Table 5: RACC-homo’s results of the comparative experiments on the length of the predefined learnable prompts $\theta_{vq}$ and $\theta_{d}$. The two sets of learnable prompts are randomly initialized here.
4.6 Broad Applicability of RACC